# GeneMAN: Generalizable Single-Image 3D Human Reconstruction from Multi-Source Human Data

### v1

@@ C.2 Quantitative Geometric Evaluation
C.3 The Ablation of Multi-Source Human Dataset
To validate the effectiveness of our constructed multi-source human dataset, we ablate each part of our multi-source human dataset: 3D scan human data, video human data, augmented human data, and 2D human data. We set the following setting: (a) “baseline” which uses the original 2D and 3D prior model (Stable Diffusion V1.5 and Zero-1-to-3) in GeneMAN, (b) “Baseline + 3D” which uses prior models only trained on 3D scanned human data in GeneMAN, (c) “Baseline + 3D + Video” which uses prior models trained on 3D scanned human data and video human data in GeneMAN, (d) “Baseline + 3D + Video + AUG” which uses prior models trained on 3D scan human data, video human data and augmented human data in GeneMAN, (e) “Ours” which use prior models trained on the complete dataset. Similar to Sec. 5.2 in the main text, we evaluate the performance of each setting using PSNR, LPIPS, and CLIP-Similarity across a set of $30$ test cases. As shown in Table 5, introducing 3D scans significantly improved the model’s consistency, resulting in a 0.046 increase in CLIP-similarity. Building on this, incorporating video data and augmented data further enhances texture quality and consistency. By utilizing the full dataset, our full-fledged method achieves the best results in both texture quality and consistency.
@@ C.3 The Ablation of Multi-Source Human Dataset
Appendix D Limitations
Although our method achieves superior reconstruction performance on in-the-wild images compared to state-of-the-art approaches, it requires a longer optimization time to generate each 3D assert compared to feed-forward methods such as PIFu [46], GTA [70], and SiTH [13], which may limit its applicability. Nevertheless, our approach is still faster than other optimization-based method TeCH [18]. The model efficiency of baseline methods and ours are shown in Fig. 6. Regarding reconstruction quality, our method still struggles to achieve fine-grained modeling for certain parts, such as the hands of full-body humans. Additionally, it lacks specific designs for handling occluded individuals. For human-with-objects reconstruction, our approach primarily focuses on reconstructing personal belongings, but it performs poorly with particularly large or complex objects, such as bicycles. Research on human-with-object reconstruction will be a key focus of our future work.
Table 6: Model Efficiency. PIFu [46], GTA [70], and SiTH [13] are feed-forward methods, whereas TeCH [18] and GeneMAN (ours) are optimization-based methods. Our GeneMAN takes 2.35 hours in total to generate each 3D human assert(Geometry Initialization takes 22 min, Geometry Sculpting takes 65 min, Latent Space Texture Refinement takes 37 min, and Pixel Space Texture Refinement takes 11 min). All methods were tested on a single NVIDIA A100 80GB GPU.
@@ Appendix D Limitations
Inference Time 4.6s 24s 3.5h 43min 2.35h

### v2

@@ D.2 Quantitative Geometric Evaluation
D.3 Ablation of Multi-Source Human Dataset
To validate the effectiveness of our constructed multi-source human dataset, we perform an ablation study on each component: 3D scan human data, video human data, augmented human data, and 2D human data. Specifically, we compare the reconstruction results using prior models trained with various combinations of these data components. The following experimental settings are evaluated: (a) “Baseline”, which replaces 2D and 3D prior models in GeneMAN with their original counterparts (Stable Diffusion V1.5 [44] and Zero-1-to-3 [28]); (b) “Baseline + 3D”, which uses prior models trained solely on 3D scanned human data; (c) “Baseline + 3D + Video”, which employs prior models trained on both 3D scanned human data and video human data; (d) “Baseline + 3D + Video + AUG” which incorporates prior models trained on 3D scan human data, video human data and augmented human data; (e) “Ours”, which utilizes prior models trained on the complete dataset. As detailed in Sec. 5.2 of the main text, we evaluate the performance of each setting using PSNR, LPIPS, and CLIP-Similarity across a set of $30$ test cases. As shown in Tab. 4, incorporating 3D scans significantly improves the model’s multi-view consistency, resulting in a 0.046 increase in CLIP-similarity. Furthermore, adding video data and augmented data enhances both texture quality and consistency. By utilizing the full dataset, our full-fledged method achieves the best performance in both texture quality and consistency.
@@ D.3 Ablation of Multi-Source Human Dataset
Table 5: Model Efficiency. Comparison of inference times between our model and the baselines. Note that PIFu [47], GTA [71], and SiTH [13] are feed-forward methods, whereas TeCH [18] and GeneMAN (ours) are optimization-based approaches that require per-subject optimization. Our GeneMAN model requires a total of 1.42 hours to generate a 3D human asset, with the following time breakdown: Geometry Initialization (22 min), Geometry Sculpting (15 min), Latent Space Texture Refinement (37 min), and Pixel Space Texture Refinement (11 min). All methods are tested on a single NVIDIA A100 80GB GPU.
@@
Inference Time 4.6s 24s 3.5h 2min 1.42h
Appendix E Limitations
Although our method achieves superior reconstruction performance on in-the-wild images compared to state-of-the-art approaches, it requires a longer optimization time to generate each 3D asset compared to feed-forward methods such as PIFu [47], GTA [71], and SiTH [13], which may limit its practical applicability. Nevertheless, our approach remains faster than TeCH [18], offering a $59.4\%$ increase in efficiency while delivering better quality. The model efficiency of both baseline methods and ours are reported in Tab. 5. Regarding reconstruction quality, our method still struggles to achieve fine-grained modeling for certain parts, such as the hands of full-body humans. Additionally, it lacks specific designs for handling occluded individuals. For human-with-objects reconstruction, our approach primarily focuses on reconstructing personal belongings, but it performs poorly with particularly large or complex objects, such as bicycles. Research on human-with-object reconstruction will be a key focus of our future work.
Figure 15: More GeneMAN Results with Complex Poses on in-the-wild Images. Best view with zoomed in.
Figure 16: More GeneMAN Results with Complex Poses on CAPE [33]. Best view with zoomed in.